# A State-Specific Iterative Decoupling Scheme Based on Perturbation Theory for Low-Energy Electronic States

### v1

@@ 2 State-specific RSBW scheme
This section details the key steps of SS-RSBW procedure. The method builds on the foundations of RS and BW perturbation theories, combining iterative effective Hamiltonian constructions with energy corrections tailored to individual electronic states. A key aspect of the approach is the progressive decoupling of selected zeroth-order states from a strategically chosen subset of the remaining zeroth-order states, enabling state-specific and accurate energy predictions while maintaining compact zeroth-order representations. The methodology is introduced through the sequential treatment of the ground and excited states, starting from a reference electronic configurations basis which is gradually refined via a sequence of effective model space optimizations and perturbative corrections.
@@ 2.3 Brillouin–Wigner energy correction
After $n_{0}^{\rm RS}$ RS steps, the zeroth-order ground-state energy is then refined using a second-order Brillouin–Wigner (BW) perturbation correction. The reference wavefunction is $|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$ and the zeroth-order energy given by $E_{0}^{(n_{0}^{\rm RS})}=\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{H}_{0}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$. Assuming that the exact BW energy expansion can be approximated by its second-order value $E_{0}^{\mbox{\tiny RSBW}}$, the BW energy correction reads[32]:
$E_{0}^{\mbox{\tiny RSBW}}$ $=E_{0}^{(n_{0}^{\rm RS})}+\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq 0}\frac{\left|\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{j}^{(n_{0}^{\rm RS})}\rangle\right|^{2}}{E_{0}^{\mbox{\tiny RSBW}}-E_{j}^{(n_{0}^{\rm RS})}}.$ (5)
This equation is either solved iteratively, or by setting $E_{0}^{\mbox{\tiny RSBW}}\approx E_{0}^{(n_{0}^{\rm RS})}$ in the energy denominators.
@@ 2.4 Iterative extension to excited states
Algorithm 1 Multi-step RSBW procedure for computing the energies of $M$ targeted states. In practice, the procedure allows to generate the lowest states starting from the Hartree-Fock $\hat{H}^{(0)}$ Hamiltonian.
@@ 4.2 Energy evaluation via BW correction
The energies of the singlet states were computed by applying second-order Brillouin-Wigner (BW) corrections to the zeroth-order energies of the optimized states $\{|\psi_{k}^{(m_{k}^{\rm RS})}\rangle\}$. The corrected energy $E_{k}^{\mbox{\tiny RSBW}}$ is defined by:
$E_{k}^{\mbox{\tiny RSBW}}$ $=E_{k}^{(m_{k}^{\rm RS})}+\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{k}^{(m_{k}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq k}\frac{\left|\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{j}^{(m_{k}^{\rm RS})}\rangle\right|^{2}}{E_{k}^{\mbox{\tiny RSBW}}-E_{j}^{(m_{k}^{\rm RS})}}.$ (7)
@@
Figure 7: SS-RSBW energies $\{E^{\rm RSBW}_{0\text{–}2}\}$ of the three lowest singlet states of LiH as a function of the internuclear distance $R$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison.
The quality of the computed energy levels was assessed using the root-mean-square (RMS) error, defined as
$\Delta^{\rm RMS}=\sqrt{\frac{1}{N_{R}}\sum_{i_{R}=1}^{N_{R}}\left(E^{\rm RSBW}_{i_{R}}-E^{\rm Exact}_{i_{R}}\right)^{2}},$
where $N_{R}=51$ denotes the number of grid points used to sample the internuclear distance $R$. The RMS errors for the three computed singlet states are reported in Table 1.
Energy Level $E^{\rm RSBW}_{0}$ $E^{\rm RSBW}_{1}$ $E^{\rm RSBW}_{2}$
$\Delta^{\rm RMS}$ (Ha) $7.1\times 10^{-5}$ $8.7\times 10^{-5}$ $3.3\times 10^{-6}$
Table 1: Root-mean-square (RMS) errors for the three lowest singlet energy states of LiH computed with the SS-RSBW method. Reference values are obtained from FCI calculations. All calculations are performed using the STO-3G AOs basis set.
Table 1 shows that, for all three singlet states, the RMS errors remain below $10^{-4}$ Hartree, indicating excellent agreement between the SS-RSBW results and the exact FCI energies over the entire range of internuclear distances. The algorithm converges rapidly, requiring less than five iterations to solve Eq. (7) across all geometries. Such observation highlights the quality of the RS-optimized zeroth-order states and the overall efficiency of the proposed scheme for accurately describing low-lying singlet states in correlated electronic systems. Besides, the enhanced convergence suggests that higher-order BW energy expansion might be included, shifting away the size-consistency drawback of the BW theory.
@@ 5.2 BW-corrected energies
Figure 11: SS-RSBW energies $\{E^{\rm RSBW}_{0\text{–}2}\}$ of the ground and two lowest singlet excited states of the H4 ring as a function of $\theta$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison. The first and second excited states crossings for $\theta=66^{\circ}$ (and symmetric counterpart $\theta=114^{\circ}$) are accurately reproduced by the SS-RSBW procedure.
Energy Level $E^{\rm RSBW}_{0}$ $E^{\rm RSBW}_{1}$ $E^{\rm RSBW}_{2}$
$\Delta^{\rm RMS}$ (Ha) $3.9\times 10^{-5}$ $2\times 10^{-5} 1.1$$1\times 10^{-5}$
Table 2: Root-mean-square (RMS) errors for the three lowest singlet energy levels of H4 computed with the SS-RSBW method. References are FCI values.
Figure 11 reports the FCI energies of the ground and the two lowest singlet excited states as a function of $\theta$, together with the results obtained from the SS-RSBW method (dashed line). The latter were obtained from a second-order BW expansion and converged after less than five iterations (see Eq. 5). The overall agreement is excellent (see Table 2) across the considered angular range. For $\theta=66^{\circ}$ and $\theta=114^{\circ}$, the two excited states are degenerate, a feature which is reproduced with high accuracy by our approach.
In practice, the procedure triggers a local enrichment of the model space during the optimization of the ZO-CS for the first excited state (see Figure 10-a), thereby explicitly including the CSFs responsible for the coupling of the two states. This targeted enlargement of the model space improves the conditioning of the BW correction and enables the method to recover two numerically degenerate levels, whereas a naïve fixed-reference perturbative strategy would likely fail. This behavior has also been analyzed in previous works on this method [1].

### v2

@@ 2 State-specific RSBW scheme
This section details the key steps of the SS-RSBW procedure. The method builds on the foundations of RS and BW perturbation theories, combining iterative effective Hamiltonian constructions with energy corrections tailored to individual electronic states. A key aspect of the approach is the progressive decoupling of selected zeroth-order states from a strategically chosen subset of the remaining zeroth-order states, enabling state-specific and accurate energy predictions while maintaining compact zeroth-order representations. The methodology is introduced through the sequential treatment of the ground and excited states, starting from a reference electronic configurations basis which is gradually refined via a sequence of effective model space optimizations and perturbative corrections.
@@ 2.3 Brillouin–Wigner energy correction
After $n_{0}^{\rm RS}$ RS steps, the zeroth-order ground-state energy is then refined using a second-order Brillouin–Wigner (BW) perturbation correction to reach SS-RSBW energy values. The reference wavefunction is $|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$ and the zeroth-order energy given by $E_{0}^{(n_{0}^{\rm RS})}=\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{H}_{0}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$. Assuming that the exact BW energy expansion can be approximated by its second-order value $E_{0}^{\mbox{\tiny SS-RSBW}}$, the BW energy correction reads[32]:
$E_{0}^{\mbox{\tiny SS-RSBW}}$ $=E_{0}^{(n_{0}^{\rm RS})}+\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq 0}\frac{\left|\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{j}^{(n_{0}^{\rm RS})}\rangle\right|^{2}}{E_{0}^{\mbox{\tiny SS-RSBW}}-E_{j}^{(n_{0}^{\rm RS})}}.$ (5)
This equation is either solved iteratively, or by setting $E_{0}^{\mbox{\tiny SS-RSBW}}\approx E_{0}^{(n_{0}^{\rm RS})}$ in the energy denominators.
@@ 2.4 Iterative extension to excited states
Algorithm 1 $S$-RSBW procedure for computing the energies of $M$ targeted states. In practice, the procedure allows to generate the lowest states starting from the Hartree-Fock $\hat{H}^{(0)}$ Hamiltonian.
@@ 4.2 Energy evaluation via BW correction
The SS-RSBW energies of the singlet states were computed by applying second-order BW corrections to the zeroth-order energies of the optimized states $\{|\psi_{k}^{(m_{k}^{\rm RS})}\rangle\}$. The corrected energy $E_{k}^{\mbox{\tiny SS-RSBW}}$ is defined by:
$E_{k}^{\mbox{\tiny SS-RSBW}}$ $=E_{k}^{(m_{k}^{\rm RS})}+\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{k}^{(m_{k}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq k}\frac{\left|\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{j}^{(m_{k}^{\rm RS})}\rangle\right|^{2}}{E_{k}^{\mbox{\tiny SS-RSBW}}-E_{j}^{(m_{k}^{\rm RS})}}.$ (7)
@@
Figure 7: SS-RSBW energies $\{E^{\text{SS-RSBW}}_{0\text{–}2}\}$ of the three lowest singlet states of LiH as a function of the internuclear distance $R$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison.
The quality of the computed energy levels obtained from both RS and SS-RSBW methods, was assessed using the root-mean-square (RMS) error, defined as
$\Delta^{\text{method}}=\sqrt{\frac{1}{N_{R}}\sum_{i_{R}=1}^{N_{R}}\left(E^{\text{method}}_{i_{R}}-E^{\rm FCI}_{i_{R}}\right)^{2}},$
where the upperscript "method" is either RS or SS-RSBW, and $N_{R}=51$ denotes the number of grid points used to sample the internuclear distance $R$. The RMS errors $\Delta^{\rm RS}$ and $\Delta^{\text{SS-RSBW}}$ for the three computed singlet energy levels are reported in Table 1.
Energy Level $E_{0}$ $E_{1}$ $E_{2}$
$\Delta^{\rm RS}$ (mHa) $1.0$ $2.3$ $0.1$
$\Delta^{\text{SS-RSBW}}$ (mHa) $0.07$ $0.08$ $0.003$
Table 1: Root-mean-square errors (mHa) for the three lowest singlet energy levels of LiH. The first row reports errors from RS treatments, while the second row corresponds to the full SS-RSBW results. Reference values are obtained from FCI calculations. All computations are performed using the STO-3G AO basis set.
Table 1 suggests that the RS treatments already provide remarkably accurate zeroth-order energies over the bond-length range, reaching spectroscopic accuracy (RMS $\sim$ 1 mHa). This result underscores the robustness of the underlying partitioning and confirms that the reference states are both accurate and systematically improvable. Building on this, the full SS-RSBW scheme further reduces the errors: for all three singlet states, the RMS deviations from the exact FCI energies remain below 0.1 mHa. The algorithm converges rapidly, requiring less than five iterations to solve Eq. (7) for all geometries. This highlights both the quality of the RS-optimized zeroth-order states and the overall efficiency of the scheme for describing low-lying singlet states in correlated electronic systems. Besides, the enhanced convergence suggests that higher-order BW energy expansion might be included, shifting away the size-consistency drawback of the BW theory.
@@ 5.2 BW-corrected energies
Figure 11: SS-RSBW energies $\{E^{\text{SS-RSBW}}_{0\text{–}2}\}$ of the ground and two lowest singlet excited states of the H4 ring as a function of $\theta$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison. The first and second excited states crossings for $\theta=66^{\circ}$ (and symmetric counterpart $\theta=114^{\circ}$) are accurately reproduced by the SS-RSBW procedure.
Energy Level $E_{0}$ $E_{1}$ $E_{2}$
$\Delta^{\rm RS}$ (mHa) $0.4$ $0.02$ $0.08$
$\Delta^{\text{SS-RSBW}}$ (mHa) $0.04$ $0.01$ $0.01$
Table 2: Root-mean-square errors (mHa) for the three lowest singlet energy levels of H4. The first row reports errors from RS treatments, while the second row corresponds to the SS-RSBW results. References are FCI values.
Figure 11 shows the FCI energies of the ground state and the two lowest singlet excited states as a function of $\theta$, together with the SS-RSBW results (dashed line). The SS-RSBW energies were obtained from a second-order BW expansion and converged with less than five iterations (see Eq. 5). Overall, the agreement is excellent (see Table 2) for both SS-RSBW and RS energies across the entire angular range considered.
For $\theta=66^{\circ}$ and $\theta=114^{\circ}$, the two excited states are degenerate, a feature which is reproduced with high accuracy by our approach. In practice, the procedure triggers a local enrichment of the model space during the optimization of the ZO-CS for the first excited state (see Figure 10-a), thereby explicitly including the CSFs responsible for the coupling of the two states. This targeted enlargement of the model space improves the conditioning of the BW correction and enables the method to recover two numerically degenerate levels, whereas a naïve fixed-reference perturbative strategy would likely fail. This behavior has also been analyzed in previous works on this method [1].